OpenStudy (anonymous):

Find the area of the region bounded by: y=-x^2 + 5 y=-x+3

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

let: f(x) = -x^2 +5, h(x) = -x +3. Find the x coordinates 'a' and 'b' of the 2 intersections by solving x for f(x) = h(x). Then the area is given by: \[ \int^b_a f(x)-h(x) \cdot dx \]

OpenStudy (anonymous):

so how would i graph that since it's "y="

OpenStudy (anonymous):

on my calculator

OpenStudy (anonymous):

or would i just plug in points

OpenStudy (anonymous):

you graph y=f(x) and y = h(x), 2 functions on the same graph\

OpenStudy (anonymous):

idk if my calculator can graph that since it's "y="

OpenStudy (anonymous):

so how would i graph the two equations, would i plug in points for "x"?

OpenStudy (anonymous):

you can just graph: y=-x^2 + 5 y=-x+3

OpenStudy (anonymous):

OHH okay ty can you help me with another one? :D

OpenStudy (agent0smith):

To find the points where the graphs intersect, just set them equal to each other and solve for x: \[\large y=-x^2 + 5 \]\[\large y=-x+3\] so let \[\large -x+3=-x^2+5\] Now solve for x, to find your limits of integration.